OpenStudy (anonymous):

WILL GIVE MEDALS TO SOMEONE WHO CLEARLY AND THOROUGHLY EXPLAINS IT!!!!!!!! a number cube is rolled 450 times. the number 3 comes up 67 times. what is the theoretical probability of rolling a 3 as a fraction? what is the experimental probability of rolling a 3 as a fraction? write in simplest form.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

ill give a medal when everyone is done answering

OpenStudy (sidsiddhartha):

a cube has 6 sides so every side can come so theoritical prob=1/6 and exp probability is simply=67/450

OpenStudy (sidsiddhartha):

u dont understand?

OpenStudy (anonymous):

i still need a bit more answers can you help @admistre64

OpenStudy (anonymous):

yes but ineed the other part of the question also but thanks for the help

OpenStudy (anonymous):

@amistre64 can you also help

OpenStudy (anonymous):

@mathmale can you help

OpenStudy (anonymous):

oh nevermind thanks